# Blinded by Generated Contexts: How Language Models Merge Generated and Retrieved Contexts for Open-Domain QA?

### v2

@@ 4.1 LLMs Prefer Self-Generated Contexts
We begin our analysis by examining LLMs’ QA performance on context-conflicting datasets to reveal how well can LLMs utilize both types of contexts. Table 2 presents the Exact Match scores (Yu et al., 2022) across various LLMs. Surprisingly, LLMs demonstrate significantly low performance ($\leq 22.16\%$) on AIR subsets, despite the fact that the retrieved context alone consistently yields the correct answer on these subsets. In contrast, most LLMs exhibit strong performance on AIG subsets (near $90\%$ or more). An exception is the Llama2-7B model, which performs relatively poorly, likely due to its inferior instruction-following capabilities. Overall, all LLMs exhibit a significant performance gap between AIR and AIG datasets, with a pronounced decline in performance when the correct answers come from retrieved contexts.
To further reveal LLMs’ behavior underlying the QA performance, we trace the source contexts of LLMs’ answers using the proposed $\mathrm{DiffGR}$ metric. An ideal LLM that optimally leverages information from both sides should always rely on retrieved contexts in case of AIR (with $\mathrm{DiffGR=-1}$), and on generated contexts for AIG (with $\mathrm{DiffGR=1}$). Contrary to expectations, Figure 5 illustrates that LLMs fail to identify the correct information and consistently tend to rely on generated contexts on both AIG and AIR subsets. This result indicates a pronounced bias in LLMs to favor generated contexts, even when they provide incorrect information. This bias leads to the insufficient utilization of retrieved contexts mentioned above, and highlights a critical challenge for existing LLMs in effectively merging generated and retrieved contexts.
As the bias on AIR subsets has more direct impact on the performance, the following experiments and analysis will focus on the biases on these subsets to conserve space. Results on the AIG subsets can be found in Appendix A.9.
@@ 4.2 LLMs Broadly Prefer Generated Contexts
Above experiments reveal the bias in LLMs to favor the self-generated context. Consequently, an intriguing question emerges: Are LLMs biased exclusively towards self-generated contexts, or do they exhibit a similar preference for contexts generated by other LLMs? To investigate this question, this section extends the experiments to more flexible combinations of generators and readers. This setting is also of practical significance, as recent works has explored the decoupling of generators and readers to achieve modularization of knowledge (Luo et al., 2023; Feng et al., 2023).
We construct context-conflicting datasets for each (generator, reader) pair respectively, as detailed in Appendix A.7. Base on these datasets, we then compute $\mathrm{DiffGR}$ metric to examine biases across various (generator, reader) pairs, as shown in Figure 6, and observe two notable insights: (i) LLMs also biased towards contexts generated by other LLMs.This suggests that such bias in LLMs is widespread and not limited to self-generated contexts. (ii) LLMs exhibit a stronger preference for contexts generated by themselves.The sole exception is Llama2-7b, which shows the strongest preference when paired with Llama2-13b as its generator. This phenomenon likely results from their highly similar model structures and training processes (Touvron et al., 2023).
@@ 5 Why LLMs Prefer Generated Contexts
In this section, we investigate why LLMs prefer generated contexts rather than retrieved contexts, from several perspectives: confirmation bias in Section 5.1, context similarity to the question in Section 5.2, and context completeness in Section 5.3.
@@ 5.1 Effect of Confirmation Bias
Recently, Xie et al. (2023) identified a confirmation bias in LLMs, wherein they exhibit a preference for contexts consistent with their parametric knowledge (or parametric memory), when faced with two conflicting generated contexts. In our experiments discussed in Section 4.1, which involve a single LLM serving both as reader and generator, it is natural to assume that the generated contexts align with the LLM’s parametric knowledge. This assumption leads to a key question: Does the confirmation bias lead to the observed preference for generated contexts?
To investigate the effect of confirmation bias, we designed controlled experiments that disrupt the consistency between generated contexts and the LLMs’ parametric knowledge. Overall, we enforce LLMs to make up a counter-memory context $d^{\varrho^{\prime}}_{1}$, which supports a same-type yet different answer compared to the original generated context. Then, we replace the generated context $d^{\varrho}_{1}$ with the counter-memory context $d^{\varrho^{\prime}}_{1}$ and recalculate $\mathrm{DiffGR}$ to assess shifts in preference after excluding confirmation bias. Table 3 reveals that LLMs maintain a significant preference for generated contexts when they contain information inconsistent with parametric knowledge. This indicates that confirmation bias does not play a major role in the preferences for generated contexts. Instead, other distinct differences between generated and retrieved contexts may be more influential, potentially overshadowing the effects of confirmation bias. In Section A.10, we further discuss the relationship between our conclusions and confirmation bias, and provide a detailed description of the experimental design process.
@@
Table 3: $\mathrm{DiffGR}$ of different input context pairs. “Gen”, “Ret” and “Ctr” respectively represent generated contexts, retrieved contexts and counter-memory contexts.
@@ 5.2 Effect of Text Similarity
The text similarity between a context and a question can reflect the degree of their relevance. To investigate the potentially effect of the similarity, we employ Jaccard similarity and BERTScore (Zhang et al., 2020) to analyze the contexts on the constructed context-conflicting datasets with the reader and generator sharing a single LLM. Figure 7 shows that generated contexts exhibit a significant higher similarity to question, even on AIR subsets where generated contexts are incorrect. This similarity discrepancy persists whether assessed by term-based overlap ($0.37$ vs. $0.18$ on TQA-AIR) or semantic similarity ($0.90$ vs. $0.86$).
Figure 7: Context-question similarity distribution of generated and retrieved contextzs on the union of AIR subsets for different LLMs (GPT-4, GPT-3.5, Llama2-13b, Llama2-7b). Distribution on AIG subsets is presented in Appendix A.11, with similar results.
Figure 8: The $\mathrm{DiffGR}$ in slices with different average $\Delta\operatorname{sim}$, with results for other LLMs in Appendix A.11.
To further clarify the influence of the observed similarity discrepancies, we rank the samples according to the similarity gap $\Delta\operatorname{sim}$ between generated and retrieved contexts, and divide the dataset into $n$ ($n=5$ here [^1] Similar results and observations are found with other $n$.) slices with an equal number of samples.
@@
Here, $\operatorname{sim}(q,d^{\varrho})$ is the similarity between generated context and question, and $\operatorname{sim}(q,d^{\gamma})$ is for retrieved context. Figure 8 illustrates the relationship between the average $\Delta\operatorname{sim}$ within each slice and the corresponding $\mathrm{DiffGR}$. From the results, we observe a general trend that on slices with a larger average similarity gap, LLMs exhibit a increased preference for generated contexts. Overall, the notable variations in $\mathrm{DiffGR}$ with respect to the similarity gap indicate that text similarity is a significant factor in the preference for generated contexts. These findings suggest that generated contexts should be applied with greater caution to mitigate the influence of highly relevant but misleading information.
To facilitate understanding why the similarity affects LLMs’ preference, we include some examples in Appendix A.11.4. From these cases, we observe that contexts with greater similarity often support candidate answers in a more straightforward manner, for instance, by mirroring the phrasing used in the questions. Conversely, the contexts with low similarity introduce more challenges, often necessitating an understanding of synonyms and even some inferences.
@@ 5.3 Effect of Context Completeness
In all above experiments, there is a key difference between generated and retrieved contexts that may affects the context preference: semantic and sentence completeness. Concretely, current retrieval systems typically employ fixed-length truncation to divide a complete article into multiple passages, which serve as the fundamental units for retrieval tasks (Karpukhin et al., 2020; Wang et al., 2019; Zhu et al., 2021). This truncation often results in retrieved contexts with incomplete semantic meaning, as well as sentences that are cut off at beginnings or endings. In contrast, generated contexts in above experiments are naturally produced by LLMs (Nature), resulting in enhanced semantic and sentence completeness.
Context Length Completeness Similarity Semantic Sentence Jaccard BERTScore Retrieved 107.4 0.114 0.801 Nature 109.7 0.202 0.879 Trunc 107.4 0.196 0.877 S-Trunc 105.9 0.193 0.876
Table 4: Contexts Statistics on NQ datasets. Appendix A.5 shows statistics on TQA, with similar conclusion.
Context Pair NQ-AIR TQA-AIR GPT-3.5 Llama2-13b GPT-3.5 Llama2-13b Nature vs. Ret 0.6562 0.5987 0.6101 0.5799 Trunc vs. Ret 0.2536 0.0958 0.3581 0.2042 S-Trunc vs. Ret 0.2394 0.1071 0.3302 0.1958
@@
To investigate the potential effects of completeness on the observed bias, we conduct controlled experiments that vary the semantic and sentence completeness of generated contexts [^2] We also attempted to control the completeness of retrieved contexts, but it was challenging to isolate it from confounding factors like length. This aspect is left for future work. using following methods: (a) Truncation (Trunc.) eliminates the length constraints from the generation prompt of Section 3, allowing LLMs to generate extended contexts. These generated contexts are then truncated to match the length of retrieved contexts, thereby simulating both semantic and sentence incompleteness of retrieved contexts. (b) Sentence Truncation (S-Trunc.): Based on method (a), we truncate generated contexts only at the end of a sentence to preserve the sentence completeness, while simulating the semantic incompleteness.
@@
We evaluate LLMs’ preference between generated versus retrieval context, varying the completeness of generated context, following the same pipeline in Section 4.1. Table 5 presents the $\mathrm{DiffGR}$ with different semantic and sentence completeness in generated contexts. A comparison between “Trunc.” and “S-Trunc.” reveals that sentence completeness does not significantly affect LLMs’ preference for generated contexts. In contrast, comparing “Nature” and “S-Trunc.”, we find a significant increase in preference for generated contexts, when they are semantically more complete. These findings indicate that LLMs are sensitive to the semantic completeness of the contexts, underscoring the necessity to investigate improved passage segmentation methods that maintain semantic completeness for current retrieval-augmented LMs.
@@ 6 Conclusion and Future work
In this study, we proposed a framework to investigate the underlying mechanisms by which LLMs merge retrieved and generated contexts. Our results reveal a pronounced bias towards generated contexts in several state-of-the-art LLMs (GPT 3.5/4, Llama2-7b/13b). We further identify two key factors contributing to this bias: higher similarity between generated contexts and questions, as well as the semantic incompleteness of retrieved contexts.
@@
7 Limitation
This study investigates the context preference of LLMs when utilizing generated and retrieved contexts, with the aim of informing future work on retrieval-augmented and generation-augmented methods. Our work has the following limitations:
  • More Task: Our experiments and analyses have been conducted open-domain question answering, a classical knowledge-intensive task. It is unclear whether LLMs show similar preference in other tasks.
  • Solution: Since this study focus on analyzing the phenomena and causes of context preference, it does not propose specific solutions to effectively mitigate this preference. Future efforts will explore various methods to alleviate such problem.
  • More Complex Conflict Scenarios: As this work preliminarily focus on the conflicts between generated and retrieved contexts, we employ one instance for each types of contexts. When multiple contexts exist from a single source, conflicts may also arise among contexts from the same sources. How LLMs manage such complex conflicts need further investigation.
@@ A.1 Related work
A.2 Generation-Augmented Approaches
@@
Generation-augmented approaches prompt LLMs to generate intermediate contexts for the final response, thereby leveraging their extensive parametric knowledge acquired during the pre-training phase on vast text corpora (Roberts et al., 2020; Petroni et al., 2019). Recent studies have demonstrated the effectiveness of these methods across a range of tasks. Liu et al. (2022) represents an early exploration into augmenting LLMs in commonsense reasoning tasks using knowledge generated by LLMs themselves. Sun et al. (2023); Yu et al. (2022) employ LLMs to produce background documents (or recitations) and subsequently use these documents to enhance LLM performance in knowledge-intensive tasks. Another researches, known as the chain-of-thought (Wei et al., 2022; Kojima et al., 2022), prompts LLMs to generate intermediate reasoning steps to enhance LLMs’ reasoning abilities. Desipte the effectiveness of these methods, the LLM-generated knowledge may contain hallucinations (Ji et al., 2023) due to LLMs’ outdated memory (De Cao et al., 2021) and limited memory for long-tail knowledge (Kandpal et al., 2023). The LLM-generated inaccuracy information could potentially mislead current retrieval model (Dai et al., 2023) and open-domain question answering systems (Pan et al., 2023).
A.3 Retrieval-Augmented Approaches

### v4

@@ 4.1 LLMs Prefer Self-Generated Contexts
We begin our analysis by examining LLMs’ QA performance on context-conflicting datasets to reveal how well can LLMs utilize both types of contexts. Table 2 presents the Exact Match scores (Yu et al., 2022) across various LLMs. Surprisingly, LLMs demonstrate significantly low performance ($\leq 22.16\%$) on AIR subsets, despite the fact that the retrieved context alone consistently yields the correct answer on these subsets. In contrast, LLMs exhibit strong performance on AIG subsets (most near $90\%$). Overall, all LLMs exhibit a significant performance gap between AIR and AIG datasets, with a pronounced decline in performance when the correct answers come from retrieved contexts.
@@
To further reveal LLMs’ behavior underlying the QA performance, we trace the source contexts of LLMs’ answers using the proposed $\mathrm{DiffGR}$ metric. An ideal LLM should always rely on retrieved contexts on AIR subsets ($\mathrm{DiffGR=-1}$), and always rely on generated contexts on AIG subsets ($\mathrm{DiffGR=1}$). Contrary to expectations, Figure 5 illustrates that LLMs fail to identify the correct information and consistently tend to rely on generated contexts on both AIG and AIR subsets. This result indicates a pronounced bias in LLMs to favor generated contexts, even when they provide incorrect information. This bias leads to the insufficient utilization of retrieved contexts mentioned above and highlights a critical challenge for existing LLMs in effectively merging generated and retrieved contexts. As the bias on AIR subsets has a more direct impact on the performance, the following experiments and analysis will focus on the biases on these subsets to conserve space. Results on the AIG subsets can be found in Appendix A.5.
@@ 4.2 LLMs Broadly Prefer Generated Contexts
The above experiments reveal the bias in LLMs to favor the self-generated context. Consequently, an intriguing question emerges: Do they also exhibit a similar bias towards contexts generated by other LLMs? To investigate this question, this section extends the experiments to more flexible combinations of generators and readers. This setting is also of practical significance, as recent works have explored the decoupling of generators and readers to achieve modularization of knowledge (Luo et al., 2023; Feng et al., 2023).
We construct context-conflicting datasets for each (generator, reader) pair respectively, as detailed in Appendix A.3. Based on these datasets, we then compute $\mathrm{DiffGR}$ metric to examine biases across various (generator, reader) pairs, as shown in Figure 6, and observe two notable insights: (i) LLMs also biased towards contexts generated by other LLMs.This suggests that such bias in LLMs is widespread and not limited to self-generated contexts. (ii) LLMs usually exhibit a stronger bias to contexts generated by themselves.The sole exception is Llama2-7b, which shows the strongest bias when paired with Llama2-13b as its generator. This phenomenon likely results from their highly similar model structures and training processes (Touvron et al., 2023).
4.3 Consistent Preferences Across Retrievers
The experiments conducted above have established the presence of the bias in LLMs against contexts retrieved by Contriever, a dense retrieval model. This section further explores whether such bias remains consistent across different retrieval models. To this end, we incorporate BM25 as a representative of sparse retrieval techniques, along with human-labeled golden passages, to mimic an ideal retrieval model (Gold). For each retrieval model, we respectively construct AIR context-conflicting datasets based on the NQ development sets from Karpukhin et al. (2020), which include the golden passages. As illustrated in Table 3, the results indicate that LLMs consistently display a pronounced bias in favor of generated contexts regardless of the retrieval model used.
Retriever Generator & Reader $\mathrm{DiffGR}$ BM25 Llama2-7b 0.5070 Contriever Llama2-7b 0.5016 Gold Llama2-7b 0.4656
Table 3: $\mathrm{DiffGR}$ of different retrievers on respectively AIR datasets constructed from NQ dev set.
@@ 5 Why LLMs Prefer Generated Contexts
In this section, we investigate why LLMs prefer generated contexts rather than retrieved contexts, from several perspectives: confirmation bias in §5.1, context similarity to the question in §5.2, and context completeness in §5.3.
@@ 5.1 Effect of Confirmation Bias
Recently, Xie et al. (2023) identified a confirmation bias in LLMs, wherein they exhibit a preference for contexts consistent with their parametric knowledge (also known as parametric memory) when faced with two conflicting generated contexts. In our experiments discussed in Section 4.1, which involve a single LLM serving both as reader and generator, it is natural to assume that the generated contexts align with the LLM’s parametric knowledge. This assumption leads to a key question: Does the confirmation bias lead to the observed preference for generated contexts in this work?
To investigate the effect of confirmation bias, we designed controlled experiments that disrupt the consistency between generated contexts and LLMs’ parametric knowledge. Overall, we enforce LLMs to make up a counter-memory context $d^{\varrho^{\prime}}_{1}$, which supports a same-type yet different answer compared to the original generated context. Then, we replace the generated context $d^{\varrho}_{1}$ with the counter-memory context $d^{\varrho^{\prime}}_{1}$ and recalculate $\mathrm{DiffGR}$ to assess shifts in preference after excluding confirmation bias, as detailed in Section A.6. Table 4 reveals that LLMs maintain a significant bias to generated contexts when they are inconsistent with LLMs’ parametric knowledge, indicating that confirmation bias does not play a major role in the observed bias. This finding does not negate the presence of confirmation bias, since other factors distinguishing generated from retrieved contexts might exert a more significant influence, potentially overshadowing the effects of confirmation bias.
Notably, GPT-3.5 even shows a stronger bias towards counter-memory contexts on TQA-AIR, a phenomenon also observed in Xie et al. (2023). This is likely attributed to the higher similarity of counter-memory contexts generated by GPT-3.5, on both our dataset and the ConflictQA datasets of Xie et al. (2023), as discussed in Section A.6.2.
@@
Table 4: $\mathrm{DiffGR}$ of different input context pairs. “Gen”, “Ret” and “Ctr” respectively represent generated contexts, retrieved contexts, and counter-memory contexts.
@@ 5.2 Effect of Text Similarity
The text similarity between a context and a question can reflect the degree of their relevance. To investigate the potential effect of the similarity, we employ Jaccard similarity and BERTScore (Zhang et al., 2020) to analyze the contexts on the constructed context-conflicting datasets with the reader and generator sharing a single LLM. Figure 7 shows that generated contexts exhibit a significantly higher similarity to the question on AIR subsets, despite the fact that generated contexts are incorrect on these subsets. This similarity discrepancy between generated and retrieved contexts persists whether assessed by term-based overlap (average $0.37$ vs. $0.18$ on TQA-AIR) or semantic similarity ($0.90$ vs. $0.86$).
Figure 7: Context-question similarity distribution of generated and retrieved contexts on the union of AIR subsets for different LLMs. Distribution on AIG subsets is presented in Appendix A.7, with similar results.
Figure 8: The $\mathrm{DiffGR}$ in slices with different average $\Delta\operatorname{sim}$. Results for other LLMs are in Appendix A.7.
To further clarify the influence of the observed similarity discrepancies, we rank the samples according to the similarity gap $\Delta\operatorname{sim}$ between generated and retrieved contexts. Then, we divide the dataset into $n$ ($n=5$ here [^2] Similar results and observations are found with other $n$.) slices with an equal number of samples.
@@
Here, $\operatorname{sim}(q,d^{\varrho})$ is the similarity between generated context and question, and $\operatorname{sim}(q,d^{\gamma})$ is for retrieved context. Figure 8 illustrates the relationship between the average $\Delta\operatorname{sim}$ within each slice and the corresponding $\mathrm{DiffGR}$. From the results, we observe a general trend that LLMs exhibit an increased bias to generated contexts on slices with a larger average similarity gap, which indicates that text similarity is a significant factor in the preference for generated contexts. These findings suggest that generated contexts should be applied with greater caution to mitigate the influence of highly relevant but misleading information.
To facilitate understanding why the similarity affects LLMs’ preference, we include some examples in Appendix A.7.4. From these cases, we observe that contexts with higher similarity often support candidate answers more straightforwardly, for instance, by mirroring the phrasing used in the questions. Conversely, the contexts with low similarity introduce more challenges, often necessitating an understanding of synonyms and even some inferences.
@@ 5.3 Effect of Context Completeness
In all the above experiments, there is a key difference between generated and retrieved contexts that may affect the context preference: semantic and sentence completeness. Concretely, current retrieval systems typically employ fixed-length truncation to divide a complete article into multiple passages, which serve as the fundamental units for retrieval tasks (Karpukhin et al., 2020; Wang et al., 2019; Zhu et al., 2021). This truncation often results in retrieved contexts with incomplete semantic meaning, as well as sentences that are cut off at beginnings or endings. In contrast, generated contexts in the above experiments are naturally produced by LLMs (Nature), resulting in enhanced semantic and sentence completeness.
Context Length Completeness Similarity Semantic Sentence Jaccard BERTScore Retrieved 107.4 0.114 0.801 Nature 109.7 0.202 0.879 Trunc. 107.4 0.196 0.877 S-Trunc. 105.9 0.193 0.876
Table 5: Average length and similarity of contexts with different completeness (details in Appendix A.2.2).
Context Pair NQ-AIR TQA-AIR GPT-3.5 Llama2-13b GPT-3.5 Llama2-13b Nature vs. Ret 0.6562 0.5987 0.6101 0.5799 Trunc. vs. Ret 0.2536 0.0958 0.3581 0.2042 S-Trunc. vs. Ret 0.2394 0.1071 0.3302 0.1958
@@
To investigate the potential effects of completeness on the observed bias, we conduct controlled experiments that vary the semantic and sentence completeness of generated contexts [^3] We also attempted to control the completeness of retrieved contexts, but it was challenging to isolate it from confounding factors like length. This aspect is left for future work.using the following methods: (a) Truncation (Trunc.) eliminates the length constraints from the generation prompt of Section 3, allowing LLMs to generate extended contexts. These generated contexts are then truncated to match the length of retrieved contexts, thereby simulating both semantic and sentence incompleteness of retrieved contexts. (b) Sentence Truncation (S-Trunc.): Based on the method (a), we truncate generated contexts only at the end of a sentence to preserve the sentence completeness, while simulating the semantic incompleteness.
@@
We evaluate LLMs’ preference between generated versus retrieval context, varying the completeness of generated context, following the same pipeline in Section 4.1. Table 6 presents the $\mathrm{DiffGR}$ with different semantic and sentence completeness in generated contexts. A comparison between “Trunc.” and “S-Trunc.” reveals that sentence completeness does not significantly affect LLMs’ preference for generated contexts. In contrast, comparing “Nature” and “S-Trunc.”, we find a significant increase in bias towards generated contexts, when they are semantically more complete. These findings indicate that LLMs tend to favor contexts with enhanced semantic completeness, underscoring the necessity to investigate improved passage segmentation methods that maintain semantic completeness for current retrieval-augmented LMs.
@@ 6 Conclusion and Future work
In this study, we propose a framework to investigate the underlying mechanisms by which LLMs merge retrieved and generated contexts. Our results reveal a pronounced bias towards generated contexts in several LLMs (GPT 3.5/4 and Llama2-7b/13b). We further identify two key factors that may contribute to this bias: higher similarity between generated contexts and questions, and the semantic incompleteness of retrieved contexts.
@@
Limitation
Our work has the following limitations:
  • This study is confined to open-domain question answering, a representative knowledge-intensive task. The behavior of LLMs across a broader spectrum of natural language processing tasks remains to be further explored.
  • This work does not propose specific solutions to effectively mitigate the observed bias, as we focus on revealing the phenomena and analyzing the causes.
  • To create a controlled environment conducive to analysis, we utilize a single instance for each context type. LLMs face increasingly intricate conflict scenarios when handling multiple contexts from each type. These conflicts emerge not only between retrieved and internally generated contexts but also among the various contexts originating from the same source (Chen et al., 2022; Xie et al., 2023).
Ethics Statement
Data
All data used in this study are publicly available and do not pose any privacy concerns.
AI Writing Assistance
In our study, we only employed ChatGPT to polish our textual expressions rather than to generate new ideas or suggestions.
@@ A.1 Related work
A.1.1 Generation-Augmented Approaches
@@
Generation-augmented approaches prompt LLMs to generate intermediate contexts for the final response, thereby leveraging their extensive parametric knowledge acquired during the pre-training phase on vast text corpora (Roberts et al., 2020; Petroni et al., 2019). Recent studies have demonstrated the effectiveness of these methods across a range of tasks. Liu et al. (2022) represents an early exploration into augmenting LLMs in commonsense reasoning tasks using knowledge generated by LLMs themselves. Sun et al. (2023); Yu et al. (2022) employ LLMs to produce background documents (or recitations) and subsequently use these documents to enhance LLM performance in knowledge-intensive tasks. Another line of research, known as the chain-of-thought (Wei et al., 2022; Kojima et al., 2022), prompts LLMs to generate intermediate reasoning steps to enhance LLMs’ reasoning abilities. Despite the effectiveness of these methods, the LLM-generated knowledge may contain hallucinations (Chen et al., 2023; Ji et al., 2023) due to LLMs’ outdated memory (De Cao et al., 2021) and limited memory for long-tail knowledge (Kandpal et al., 2023). The LLM-generated inaccuracy information could potentially mislead current retrieval model (Dai et al., 2023) and open-domain question answering systems (Pan et al., 2023).
A.1.2 Retrieval-Augmented Approaches